OpenStudy (anonymous):

An angle in standard position measures 5π/6 radians and intercepts a circle with center at (0,0) and radius 4. Find the length of the arc determined by the points of intersection. HELP PLEASE :(

OpenStudy (anonymous):

So how much around a circle is 5pi/6?

OpenStudy (jdoe0001):

well let's see

OpenStudy (jdoe0001):

5/6 of a \(\pi\) that is  so, we know the angle, "in radians" we also know the radius of the circle it intercepts so, how long is the "arc"? well \(\large \textit{arc's length}=s=r\theta\qquad \begin{cases} r=radius\\ \theta=\textit{angle in radians} \end{cases} \)